# A Review of Barren Plateaus in Variational Quantum Computing

### v1

@@ I Introduction
Variational quantum computing has gathered momentum in recent years as a promising approach for quantum computers [1, 2, 3, 4, 5, 6, 7]. At their core, most variational quantum algorithms follow the same principle: Convert a problem into an optimization task, and use a classical device to train a parametrized quantum learning model. The popularity of this computational paradigm can be attributed in part to its flexibility, opening up applications in diverse areas in basic science and machine learning, but also because of the hope that it is more robust to the constraints of near-term quantum hardware.
In contrast to conventional quantum algorithms, which generally come with provable guarantees, variational quantum algorithms share much of the heuristic nature of classical machine learning. However, as opposed to classical machine learning, we do not yet have the hardware required to test these algorithms at scale. Moreover, it quickly becomes clear to any practitioner that out-of-the-box variational schemes can be hard to train as the number of qubits becomes larger and larger. This has prompted efforts to analyze the scalability of these algorithms and to determine trainability barriers that could obstruct their practical use in realistic problem sizes.
@@ II Variational Quantum Computing
Figure 1: Variational quantum computing. An $n$-qubit quantum computer is initialized to a state $\rho$ that is evolved through a PQC $\mathcal{U}_{\bm{\theta}}$. At the output of the circuit we perform a finite set of measurements which are used to estimate a loss function $\ell_{\bm{\theta}}(\rho,O)$ as in Eq. (3). A classical computer takes as input the finite sample estimate of $\ell_{\bm{\theta}}(\rho,O)$ (or its partial derivatives) and attempts to find new sets of parameters which train the PQC and minimize the loss. As such, at the core of variational quantum computing lies the manipulation and comparison—via polynomially many samples—of the exponentially large operators $\rho(\bm{\theta})=\mathcal{U}_{\bm{\theta}}(\rho)$ and $O$.
@@ II.1 Framework
The key idea behind variational quantum computing is to combine the power of both quantum and classical computers. In this hybrid computational paradigm, one uses the exponentially large Hilbert space of the quantum device to process information and estimate some quantity of interest, and then feeds those outputs to a classical computer that analyzes them and (potentially) determines the next experiment on the quantum computer. Many such hybrid algorithms employ an optimization, or a learning-based, approach whereby the quantum experiments are characterized by a sequence of parameters that a classical optimizer attempts to train in order to solve a given task. Given that variational computing reduces the burden on quantum hardware, they are naturally well suited for near-term devices. However, their usage is by no means confined to the near-term and it is hoped that variational schemes will further flourish in the fault-tolerant era when noise is no longer an issue.
It has become standard to divide variational quantum computing schemes into two main categories: Variational Quantum Algorithms (VQAs) [1, 2, 3] and Quantum Machine Learning (QML) models [4, 5, 6]. VQAs can be considered as problem-driven, since the algorithms tackle a specific task such as preparing the ground state of some Hamiltonian. QML models on the other hand are data-driven. QML has been envisioned for virtually all learning tasks such as supervised and unsupervised learning, generative modeling, etc [6]. Throughout this work, we will study in parallel VQAs and QML models as they mostly share a common structure.
@@
As shown in Fig. 1, in variational quantum computing one first initializes the quantum system to some $n$-qubit initial state $\rho$, belonging to $\mathcal{B}(\mathcal{H}_{0})$, the set of bounded linear operators on an $2^{n}$-dimensional complex Hilbert space $\mathcal{H}_{0}$. In VQAs, $\rho$ is usually some easy-to-prepare fiducial state, while in QML $\rho$ is taken from a set (or distribution) of training states that encode either quantum or classical data (see Box II.2).
[t] Box 1: Quantum and classical data for QML By quantum data, we mean that $\rho$ is obtained from some quantum mechanical process of interest within a quantum device. By classical data $\bm{x}\in\mathcal{X}$ we mean data from some domain $\mathcal{X}$ that can be efficiently stored in a classical computer (e.g., images, financial information, events from a particle accelerator), and that we want to process in a quantum computer. Such classical data needs to be encoded in a quantum state through an embedding map $\mathcal{E}:\mathcal{X}\rightarrow\mathcal{B}(\mathcal{H}_{0})$. We also note that the classical data $\bm{x}$ can also be used to construct the PQC in Eq. (1) through data re-uploading methods [12].
@@
where $\bm{\theta}=(\bm{\theta}_{1},\ldots,\bm{\theta}_{L})$ is a set of trainable parameters. In the presence of hardware noise, or if the parameterized quantum circuit includes adding or discarding qubits, the PQC can be more generally expressed as a sequence of $L$ parametrized channels
@@
[t] Box 2: Inductive biases Inductive biases, broadly speaking, capture the assumptions and choices we make when constructing the quantum algorithm. In the pipeline of variational quantum computing these come in many forms. For instance, when working with VQAs, while there is a significant amount of freedom when choosing the input state, the measurement operator is usually more restricted. This is due to the fact that in most VQAs the problem information is encoded in the measurement operator (e.g., in the variational quantum eigensolver [13] we measure the expectation value of the Hermitian operator whose ground state we want to prepare). In QML, the input states tend to be fixed, with the encoding scheme (for classical data), the measurement operator and choice of loss function being generally more arbitrary, and thus leading to extra design freedom. Similarly, while Eq. (1) shows a general PQC, different instantiations of these computational models are distinguished form each other by the inductive biases that go into $\mathcal{U}_{\bm{\theta}_{l}}$, i.e., the choices in their specific form. This includes properties such as the number and types of gates, their placement, and how the parameters are initialized.
@@
Despite the (relative) simplicity of the variational quantum computing framework, solving Eq. (4) and training the model’s parameters can be a hard task. For instance, the optimizer might get stuck in a sub-optimal local minima [14, 15, 16, 17, 18, 19], there might not exist a solution in the parameter hyperspace [20, 21], or the landscape might exhibit BPs, the focus of this review. Importantly, and as we will see below, the hardness of training the loss will strongly depend on the inductive biases of the quantum algorithm (see Box II.2).
[t] Box 3: Effects of finite sampling In practice, we can only estimate the expectation value of the loss function $\ell_{\bm{\theta}}(\rho,O)=\mathbb{E}_{\rho(\bm{\theta})}[O]$ up to some finite precision determined by the number of measurement shots. To differentiate, $\ell_{\bm{\theta}}(\rho,O)$ from its $N$-shot approximation, we use ¯$\overline{\ell}_{\bm{\theta}}(\rho,O)$ to denote the latter. Assuming that the experiment repetition used to obtain ¯$\overline{\ell}_{\bm{\theta}}(\rho,O)$ are independent, we recover the sample variance (denoted as ${\rm Var}_{s}$) of the estimate $\begin{split}{\rm Var}_{s}[\overline{\ell}_{\bm{\theta}}(\rho,O)]&=\frac{{\rm Var}_{\rho(\bm{\theta})}[O]}{N}\\ &=\frac{\mathbb{E}_{\rho(\bm{\theta})}[O^{2}]-\mathbb{E}_{\rho(\bm{\theta})}[O]^{2}}{N}\,.\end{split}$ (5) Equation (5) quantifies the uncertainty to which ¯$\overline{\ell}_{\bm{\theta}}(\rho,O)$ is computed. Clearly, there are two contributions to this uncertainty. First, the one in the nominator arises from the fact that $\rho(\bm{\theta})$ is generally not an eigenstate of $O$, and hence different measurements can yield different outcomes. Second, the factor $1/N$ indicates that we are estimating the expectation value with a finite number of samples. Thus, in regions of a landscape with exponentially small gradients, exponentially many shots are typically required to determine a loss minimizing direction.
@@ III Types of Barren Plateaus and Loss Concentration
We say that a variational quantum computing scheme exhibits a BP if its loss function, or its gradients, concentrate exponentially about their mean in the number of qubits $n$. Intuitively, this means that the optimization landscape is mostly flat and featureless and that slightly changing the model’s parameters $\bm{\theta}$ results in only an exponentially small change in $\ell_{\bm{\theta}}(\rho,O)$ or $\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)=\partial\ell_{\bm{\theta}}(\rho,O)/\partial\theta_{\mu}$ (where $\theta_{\mu}\in\bm{\theta}$). One can readily see that this will cause serious issues when minimizing the loss function, as optimizers find descent directions by comparing the loss at different points (gradient-free) or by using derivative information (e.g., gradient-based). In particular, since we can only estimate expectation values from a quantum computer with a finite number of measurements $N$, these come with a statistical uncertainty that scales inversely proportional to $\sqrt{N}$ (see Box II.2). Thus, if we want to resolve exponentially small changes and optimize the loss, we need to spend an exponential amount of shots, rendering the algorithm inefficient and non-scalable. Indeed, if a smaller number of shots are used, the optimizer will follow changes produced by the uninformative statistical fluctuations, meaning that the parameter updates lead to a useless random walk in the loss landscape.
@@ III.1 Probabilistic concentration and narrow gorges
The most common type of loss function concentration encountered in the literature is known as probabilistic concentration. By defining $\mathbb{E}_{\bm{\theta}}$ as the expectation value over the parameter landscape, we say a loss exhibits a probabilistic BP if
@@
for some $b>1$, and for some (but not necessarily all) $\mu$. This implies through Chebyshev’s inequality that for any $\delta>0$
${\rm Pr}(|\ell_{\bm{\theta}}(\rho,O)-\mathbb{E}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]|\geqslant\delta)\in\mathcal{O}\left(\frac{1}{b^{n}}\right)\,,$ (7)
or ${\rm Pr}(|\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)-\mathbb{E}_{\bm{\theta}}[\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)]|\geqslant\delta)\in\mathcal{O}\left(\frac{1}{b^{n}}\right)$. These equations indicate that the probability that the loss, or its partial derivative, deviate by more than $\delta$ (for some $\delta\in\Omega(1/\operatorname{poly}(n))$) from its mean is exponentially small.
Here we find it important to remark that while the definition of a BP in terms of the loss is unambiguous, it is less so when stated in terms of partial derivatives. As such, our statement above that a BP arises when some (but not all) partial derivatives must concentrate highlights that concentration of the loss is a stronger means of diagnosing a BP. In fact, it is entirely possible that a loss contains some partial derivatives that are concentrated while others are not. Hence, care must be taken when establishing the absence of a BP by studying ${\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]$ as it is entirely possible that a non-concentrated loss arises only from a few well-behaved set of parameters (and thus still has a BP). However, when every partial derivative is concentrated one can prove that loss concentration and partial derivative concentration are equivalent [10, 22]. Similarly, one can show that if first-order derivatives concentrate, then so do higher-order ones [23].
Interestingly, while a probabilistic BP implies a mostly featureless, flat landscape, there can still be regions with large gradients around some minima (sometimes called fertile valleys [24]). The caveat is that these regions have to be small and the minima therein are always exponentially narrow (meaning that their relative volume in parameter space is always exponentially small [10]). These properties led researchers to name these minima as narrow gorges. Here, it is worth mentioning that while narrow gorges bring to mind an image of a very steep well, the slope of the walls of a narrow gorge cannot be too large as it bounded by the loss’s Lipschitz constant which in most cases is $\mathcal{O}(\operatorname{poly}(n))$ [10].
@@ III.2 Deterministic concentration
While probabilistic BPs lead to mostly flat and featureless landscapes, they still admit the presence of well-behaved regions such as narrow gorges. In contrast, there exists a second type of concentration, known as deterministic BPs, where all of the landscape is truly flat [25]. A deterministic BP arises when $\forall\bm{\theta}$
$|\ell_{\bm{\theta}}(\rho,O)-\mathbb{E}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]|\in\mathcal{O}\left(\frac{1}{b^{n}}\right)\,,$ (8)
for some $b>1$. Clearly, Eq. (8) implies that for all parameter values $|\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)-\mathbb{E}_{\bm{\theta}}[\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)]|\in\mathcal{O}\left(\frac{1}{b^{n}}\right)$. From here it is obvious that a deterministically concentrated loss cannot admit any type of extrema that is well separated from the loss’s mean. In fact, Eq. (8) precludes any type of non-exponentially suppressed features in the landscape.

### v2

@@ I Introduction
Variational quantum algorithms have gathered momentum in recent years as a promising computational approach [1, 2, 3, 4, 5, 6, 7]. At their core, most variational quantum algorithms follow the same principle: Convert a problem into an optimization task, and use a classical device to train a parametrized quantum learning model. The popularity of this computational paradigm can be attributed in part to its flexibility, opening up applications in diverse areas in basic science and machine learning, but also because of the hope that it is more robust to the constraints of near-term quantum hardware.
In contrast to conventional quantum algorithms, which generally come with provable guarantees, variational quantum algorithms share much of the heuristic nature of classical machine learning. Unlike classical machine learning, quantum computing lacks the hardware needed to test these algorithms at scale. Moreover, it quickly becomes clear to any practitioner that out-of-the-box variational schemes can be hard to train as the number of qubits becomes larger and larger. This has prompted efforts to analyze the scalability of these algorithms and to determine trainability barriers that could obstruct their practical use in realistic problem sizes.
@@ II Variational Quantum Computing
Figure 1: Variational quantum computing. An $n$-qubit quantum computer is initialized to a state $\rho$ that is evolved through a Parametrized Quantum Circuit (PQC) $\mathcal{U}_{\bm{\theta}}$. At the output of the circuit we perform a finite set of measurements which are used to estimate a loss function $\ell_{\bm{\theta}}(\rho,O)={\rm Tr}[\mathcal{U}_{\bm{\theta}}(\rho)O]$. A classical computer takes as input the finite sample estimate of $\ell_{\bm{\theta}}(\rho,O)$ (or its partial derivatives) and attempts to find new sets of parameters which train the PQC and minimize the loss. As such by writing the loss in the vectorized form $\ell_{\bm{\theta}}(\rho,O)=\langle\rho(\bm{\theta}),O\rangle$ one can see that at the core of variational quantum computing lies the manipulation and comparison—via polynomially many samples—of the exponentially large operators $\rho(\bm{\theta})=\mathcal{U}_{\bm{\theta}}(\rho)$ and $O$.
@@ II.1 Framework
The key idea behind variational quantum computing is to combine the power of both quantum and classical computers. This hybrid paradigm leverages the quantum device’s large Hilbert space to process information and estimate quantities of interest. The results are then fed to a classical computer, which analyzes them and determines the next quantum experiment. Many such hybrid algorithms employ an optimization or learning-based approach, where a classical optimizer trains quantum experiments characterized by a sequence of parameters to solve a task. Since variational computing reduces the burden on quantum hardware, it is naturally well-suited for near-term devices. However, their usage is not confined to the near-term, and variational schemes are expected to flourish in the fault-tolerant era when noise is no longer an issue.
[t] Box 1: Quantum and classical data for QML By quantum data, we mean that $\rho$ is obtained from some quantum mechanical process of interest within a quantum device. By classical data $\bm{x}\in\mathcal{X}$ we mean data from some domain $\mathcal{X}$ that can be efficiently stored in a classical computer (e.g., images, financial information, events from a particle accelerator), and that we want to process in a quantum computer. Such classical data needs to be encoded in a quantum state through an embedding map $\mathcal{E}:\mathcal{X}\rightarrow\mathcal{B}(\mathcal{H}_{0})$. We also note that the classical data $\bm{x}$ can also be used to construct the Parametrized Quantum Circuit (PQC) in Eq. (1) through data re-uploading methods [12].
It has become standard to divide variational quantum computing schemes into two main categories: Variational Quantum Algorithms (VQAs) [1, 2, 3] and Quantum Machine Learning (QML) models [4, 5, 6]. VQAs can be considered as problem-driven, since the algorithms tackle a specific task such as preparing the ground state of some Hamiltonian. QML models on the other hand are data-driven. QML has been envisioned for virtually all learning tasks such as supervised and unsupervised learning, generative modeling, etc [6]. Throughout this work, we will study VQAs and QML models together, as they mostly share a common structure.
@@
As shown in Fig. 1, in variational quantum computing one first initializes the quantum system to some $n$-qubit initial state $\rho$, belonging to $\mathcal{B}(\mathcal{H}_{0})$, the set of bounded linear operators on an $2^{n}$-dimensional complex Hilbert space $\mathcal{H}_{0}$. In VQAs, $\rho$ is usually an easy-to-prepare fiducial state, while in QML $\rho$ is taken from a set (or distribution) of training states that encode either quantum or classical data (see Box II.1).
@@
where $\bm{\theta}=(\bm{\theta}_{1},\ldots,\bm{\theta}_{L})$ is a set of trainable parameters. When hardware noise is present, or when the parameterized quantum circuit involves adding or removing qubits, the PQC is expressed more generally as a sequence of $L$ parametrized channels
@@
[t] Box 2: Inductive biases Inductive biases, broadly speaking, capture the assumptions and choices we make when constructing the quantum algorithm. In the pipeline of variational quantum computing these come in many forms. For instance, when working with VQAs, while there is a significant amount of freedom when choosing the input state, the measurement operator is usually more restricted. This is due to the fact that in most VQAs the problem information is encoded in the measurement operator (e.g., in the variational quantum eigensolver [13] we measure the expectation value of the Hermitian operator whose ground state we want to prepare). In QML, the input states tend to be fixed, with the encoding scheme for classical data, the measurement operator, and the choice of loss function being generally more flexible, offering additional design freedom."Similarly, while Eq. (1) shows a general PQC, different instantiations of these computational models differ based on the inductive biases that influence $\mathcal{U}_{\bm{\theta}_{l}}$, i.e., the choices in their specific form. This includes properties such as the number and types of gates, their placement, and how the parameters are initialized.
@@
Despite the (relative) simplicity of the variational quantum computing framework, solving Eq. (4) and training the model’s parameters can be a hard task. For instance, the optimizer might get stuck in a sub-optimal local minima [14, 15, 16, 17, 18, 19, 20], there might not exist a solution in the parameter hyperspace [21, 22], or the landscape might exhibit BPs, the focus of this review. Importantly, and as we will see below, the hardness of training the loss will strongly depend on the inductive biases of the quantum algorithm (see Box II.2).
[t] Box 3: Effects of finite sampling In practice, we can only estimate the expectation value of the loss function $\ell_{\bm{\theta}}(\rho,O)=\mathbb{E}_{\rho(\bm{\theta})}[O]$ up to some finite precision determined by the number of measurement shots. To differentiate, $\ell_{\bm{\theta}}(\rho,O)$ from its $N$-shot approximation, we use ¯$\overline{\ell}_{\bm{\theta}}(\rho,O)$ to denote the latter. Assuming that the experiment repetition used to obtain ¯$\overline{\ell}_{\bm{\theta}}(\rho,O)$ are independent, we recover the sample variance (denoted as ${\rm Var}_{s}$, and also known as the standard deviation of the mean [23]) of the estimate $\begin{split}{\rm Var}_{s}[\overline{\ell}_{\bm{\theta}}(\rho,O)]&=\frac{{\rm Var}_{\rho(\bm{\theta})}[O]}{N}\\ &=\frac{\mathbb{E}_{\rho(\bm{\theta})}[O^{2}]-\mathbb{E}_{\rho(\bm{\theta})}[O]^{2}}{N}\,.\end{split}$ (5) Equation (5) quantifies the uncertainty to which ¯$\overline{\ell}_{\bm{\theta}}(\rho,O)$ is computed. Clearly, there are two contributions to this uncertainty. First, the one in the numerator arises from the fact that $\rho(\bm{\theta})$ is generally not an eigenstate of $O$, and hence different measurements can yield different outcomes. Second, the factor $1/N$ indicates that we are estimating the expectation value with a finite number of samples, and therefore. Via Chebychev’s inequality this implies that the statistical fluctuations of the mean estimator are of order $1/\sqrt{N}$. Thus, in regions of a landscape with exponentially small gradients, exponentially many shots are typically required to determine a loss minimizing direction.
@@ III Types of Barren Plateaus and Loss Concentration
We define a barren plateau (BP) in a variational quantum computing scheme as occurring when the loss function, or its gradients, become exponentially concentrated around their mean as the number of qubits $n$ increases. Intuitively, this means that the optimization landscape is mostly flat and featureless and that slightly changing the model’s parameters $\bm{\theta}$ results in only an exponentially small change in $\ell_{\bm{\theta}}(\rho,O)$ or $\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)=\partial\ell_{\bm{\theta}}(\rho,O)/\partial\theta_{\mu}$ (where $\theta_{\mu}\in\bm{\theta}$). This creates significant challenges for minimizing the loss function, as optimizers determine descent directions by comparing the loss at different points. In particular, since we can only estimate expectation values from a quantum computer with a finite number of measurements $N$, these come with a statistical uncertainty that scales inversely proportional to $\sqrt{N}$ (see Box II.2). To resolve exponentially small changes and optimize the loss, we would need an exponentially large number of measurements (shots), which makes the algorithm inefficient and non-scalable. Indeed, if a smaller number of shots are used, the optimizer will follow changes produced by the uninformative statistical fluctuations, meaning that the parameter updates lead to a useless random walk in the loss landscape.
@@ III.1 Probabilistic concentration and narrow gorges
The most common type of loss function concentration encountered in the literature is known as probabilistic concentration. By defining $\mathbb{E}_{\bm{\theta}}$ as the expectation value with respect to $\bm{\theta}$ sampled from some domain according to some distribution, we say a loss exhibits a probabilistic BP if
@@
for some $b>1$, and for some but (not necessarily all) parameters $\theta_{\mu}\in\bm{\theta}$. As discussed later in Sec. IV, the choice of an ansatze $U(\bm{\theta})$, an initial state $\rho$ and an observable $O$ can individually induce the concentration. The exact value of the exponent base $b$ implicitly depends on $U(\bm{\theta})$, $\rho$ and $O$.
At this point, we note that one should understand probabilistic BPs as an average properties of a loss landscape, which means that “most” points in the landscape should be exponentially flat. To see this more formally, the exponential vanishing variance implies through Chebyshev’s inequality that for any $\delta>0$
${\rm Pr}_{\bm{\theta}}(|\ell_{\bm{\theta}}(\rho,O)-\mathbb{E}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]|\geqslant\delta)\in\mathcal{O}\left(\frac{1}{b^{n}}\right)\,,$ (7)
or ${\rm Pr}(|\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)-\mathbb{E}_{\bm{\theta}}[\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)]|\geqslant\delta)\in\mathcal{O}\left(\frac{1}{b^{n}}\right)$. These equations show that the probability of the loss or its partial derivative deviating by more than $\delta$ from its mean is exponentially small, where $\delta\in\Omega(1/\operatorname{poly}(n))$.
Here we find it important to remark that while the definition of a BP in terms of the loss is clear, it becomes less clear when expressed in terms of partial derivatives. As mentioned, a BP arises when some, but not all, partial derivatives are concentrated, indicating that analyzing the concentration of the overall loss is a more reliable way to diagnose a BP. In fact, it is entirely possible that a loss contains some partial derivatives that are concentrated while others are not. Hence, care must be taken when establishing the absence of a BP by studying ${\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]$ as it is entirely possible that a non-concentrated loss arises only from a few well-behaved set of parameters (and thus still has a BP). However, when every partial derivative is concentrated one can prove that loss concentration and partial derivative concentration are equivalent [10, 24]. Similarly, one can show that if first-order derivatives concentrate, then so do higher-order ones [25].
Although a probabilistic BP implies a largely featureless landscape, regions with large gradients around some minima (sometimes referred to as fertile valleys [26]) can still exist. The caveat is that these regions have to be small and the minima therein are always exponentially narrow (meaning that their relative volume in parameter space is always exponentially small [10]). These properties led researchers to name these minima as narrow gorges. While this term suggests a steep well, the slope of the walls in these gorges is bounded by the loss’s Lipschitz constant, which is typically in $\mathcal{O}(\operatorname{poly}(n))$ [10]. Lastly, since these are the most common BPs discussed in this review, we henceforth assume BPs to be probabilistic unless stated otherwise.
@@ III.2 Deterministic concentration
While probabilistic BPs lead to mostly flat and featureless landscapes, they still admit the presence of well-behaved regions such as narrow gorges. In contrast, there exists a second type of concentration, known as deterministic BPs, where all of the landscape is truly flat [27, 28, 29, 30]. A deterministic BP arises when the loss function’s deviation from its mean is bounded for all parameter values, $\forall\theta$, by an exponentially small term. That is, when
$|\ell_{\bm{\theta}}(\rho,O)-\ell_{0}]|\in\mathcal{O}\left(\frac{1}{b^{n}}\right)\,,$ (8)
for some $b>1$, and where $\ell_{0}$ is a concentration point which is usually independent of $\bm{\theta}$. Deterministic concentration thus arises due to the state $\rho(\bm{\theta})$ only having exponentially small overlap with $O$ for all values of $\bm{\theta}$ (which therefore implies that this BP phenomenon depends on the initial state and measurement operator). As further discussed below, this situation can arise due to high entanglement in an input state (see Sec. IV.3) and/or unital noise in PQC (see Sec. IV.4). Clearly, Eq. (8) implies that for all parameter values $|\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)-\ell_{0}|\in\mathcal{O}\left(\frac{1}{b^{n}}\right)$. A deterministically concentrated loss cannot have any extrema that are significantly separated from the loss’s mean. In fact, Eq. (8) precludes any type of non-exponentially suppressed features in the landscape.